OpenStudy (anonymous):

factor the trinomial completely r^2-14r+33

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 33 (last term) and add to -14 (middle coefficient)

OpenStudy (anonymous):

i got (r+11)(r-3)

jimthompson5910 (jim_thompson5910):

that is incorrect

jimthompson5910 (jim_thompson5910):

notice how +11*(-3) = -33 but we should have +33

OpenStudy (anonymous):

so both have to be addition

jimthompson5910 (jim_thompson5910):

the two numbers both have to be negative (since negative times negative = positive)

jimthompson5910 (jim_thompson5910):

they have to negative because the middle coefficient is something negative while the last term is positive

jimthompson5910 (jim_thompson5910):

what two negative numbers multiply to 33 ?

OpenStudy (anonymous):

11 and 3

jimthompson5910 (jim_thompson5910):

I think you mean -11 and -3

jimthompson5910 (jim_thompson5910):

so r^2-14r+33 factors to (r-11)(r-3)

jimthompson5910 (jim_thompson5910):

notice how -11*(-3) = +33 and -11 + (-3) = -14